# Interpretable Time Series Autoregression for Periodicity Quantification

### v1

@@ 7.1.1 Ridesharing Trip Time Series
7.1.2 Comparison among NNSP, MIO-DVP, and MIO
Basically, we examine three solution algorithms—NNSP, MIO-DVP, and MIO—for addressing Problem (4) on the ridesharing pickup trip time series across five years. The sparsity levels in these algorithms are set as $\tau=4,6$. In particular, we test the MIO-DVP algorithm with different sparsity levels $\tau_{0}=5,10$ for pruning decision variables, while NNSP is an important baseline for measuring the performance gain. Table II shows the objective function values of SAR with different algorithmic settings. As can be seen, the NNSP algorithm fails to find the solution as good as the MIO algorithm. The MIO-DVP algorithm performs better than NNSP if a well-designed pruning strategy such as NNSP with $\tau_{0}=10$ is available. Herein, the sparsity level $\tau_{0}$ is an influential factor in the MIO-DVP algorithm for finding a good solution as MIO. In contrast to MIO, one remarkable advantage of MIO-DVP is the efficiency and scalability due to the reduction of search space.
7.1.3 Periodicity of Ridesharing Trips
As shown in Fig. LABEL:NYC_JFK_airport_heatmap, the ridesharing pickup/dropoff trip time series is time-varying in a long-term range. In the time series, one can observe the remarkable reduction of total trip counts in 2020 due to the COVID-19 pandemic, recovering slowly until the end of 2021. To identify such kind of time-varying system behaviors, we use the proposed TV-SAR model. The time series of each month is regarded as a time segment. The order of autoregression is set as $d=168$, referring to a weekly cycle. The sparsity level is set as $\tau=4$ which aims to cover the indices of daily and weekly periodicities at $k=24$ and $k=168$, respectively.
@@ 7.1.3 Periodicity of Ridesharing Trips
Of the results in Fig. LABEL:NYC_JFK_w_tau_3, the proposed model identifies the support set as $\operatorname{supp}(\boldsymbol{w}_{\gamma})=\{1,24,167,168\},\,\forall\gamma\in[2,60]$, corresponding to local, daily, and weekly auto-correlations. Fig. LABEL:NYC_JFK_pickup_w_tau_3 shows the periodicity pattern changes of pickup trips from 2019 to 2023. As can be seen, the strength reduction of weekly periodicity and the increasing strength of daily periodicity are remarkable in 2020. The similar results are also demonstrated in Fig. LABEL:NYC_JFK_dropoff_w_tau_3 for dropoff trips, namely, the decreasing strength of weekly periodicity and the increasing strength of daily periodicity in 2020. Another finding from Fig. LABEL:NYC_JFK_dropoff_w_tau_3 is the increasing local auto-correlations at $k=1$ in 2020, implying increasing variability in the dropoff trips. Observing the coefficients at $k=168$, the weekly periodicity of dropoff trips in 2019, 2021, 2022, and 2023 are stronger than pickup trips, and such evidence is consistent with pickup and dropoff trip time series in Fig. LABEL:NYC_JFK_airport_heatmap. Although the periodicity patterns in 2020 is very different from the years before and after COVID-19, the local and nonlocal auto-correlations in 2019, 2021, 2022, and 2023 demonstrate no remarkable difference. This implies that the regularity of human mobility after COVID-19 has recovered to 2019.
7.2 Seasonality Patterns of North America Climate Variables
Daymet provides monthly climate variables (e.g., minimum/maximum temperature, and precipitation) at a spatial resolution of 1 km $\times$ 1 km [35]. [^1] https://daac.ornl.gov/DAYMET In this work, we consider the data spanning from 1980 to 2019, which is in the form of multidimensional time series. Fig. 10 shows the year-to-year variations and the temporal variability of the minimum temperature data of January in the past decade. These panels reveal seasonal temperature patterns and changes over the decade, demonstrating regional climate trends.
7.2.1 Robustness of Seasonality Patterns with Different Spatial Resolutions
Since the original dataset is overly high-resolution, we aggregate the data with different spatial resolutions, including 5 km $\times$ 5 km, 10 km $\times$ 10 km, and 20 km $\times$ 20 km. To quantify yearly seasonality of these climate time series, we use the proposed STV-SAR model with a prescribed sparsity level $\tau=3$. For instance, on the minimum temperature dataset, we build the monthly time series across $M\times N$ grids of each decade. The global support set is optimized as $\Omega=\{1,11,12\}$ with the sparsity level $\tau=3$, and the coefficients at the index $k=12$ are used to quantify the yearly seasonality of monthly time series. Fig. LABEL:fast_mip_daymet_20km_tmin_w12_sparsity3_10s shows that the strengths of yearly seasonality with different spatial resolutions are consistent, in which a relatively higher coefficient implies stronger seasonality. In 2010s, both spatial patterns of yearly seasonality highlight some very seasonal regions such as north areas of Canada, western areas of USA, and southeastern areas of USA. The temperature of central regions in North America is extremely less seasonal than other regions. As shown in Fig. LABEL:fast_mip_daymet_20km_tmin_w12_sparsity3_10s, temperature data of Mexico demonstrate remarkable variability and inconsistency among nearby areas, one can observed detailed spatial patterns of temperature seasonality from Fig. LABEL:fast_mip_daymet_5km_tmin_mexico_10s.
To address the computational challenges posed by the large-scale MIO problem in Eq.(9), we propose a two-stage optimization scheme aimed at improving scalability in the presence of thousands to millions of decision variables. TableIII summarizes the model’s running times at three spatial resolutions. As the results show, the proposed scheme enables efficient computation even for datasets with up to 3 million time series, making it well-suited for large-scale, real-world applications.
@@ 7.2.1 Robustness of Seasonality Patterns with Different Spatial Resolutions
TABLE III: Running times of STV-SAR on the North America (minimum) temperature dataset with different spatial resolutions. Note that $\nu$ denotes the number of time series.
5 km $\times$ 5 km 10 km $\times$ 10 km 20 km $\times$ 20 km
@@
Cost 48.55 s 11.77 s 2.88 s
7.2.2 Variations of Seasonality Patterns across Four Decades
Discovering the variations of seasonality patterns allows one to quantify the evolution of dynamical climate systems. Fig. LABEL:daymet_w12_sparsity3 shows yearly seasonality of monthly time series data in North America across past four decades with a spatial resolution of 10 km $\times$ 10 km. Recall that higher coefficients in the panels indicate regions with stronger yearly seasonality. We first summarize the findings from the seasonality patterns of minimum and maximum temperature data. Over the decades, regions with stronger yearly seasonality are predominantly located in northern and high-latitude areas such as Canada and Alaska. The southern regions such as Mexico seem to be less seasonal in the past four decades than northern regions, and the variation of yearly seasonality in Mexico is remarkable among relatively small areas. A remarkable expansion of highly seasonal regions occurs in 2000s compared to earlier decades, demonstrating an intensification of seasonal temperature variations. Such kind of expansion of regions with a strong yearly seasonality highlights area where temperature-driven seasonal cycles are more predictable. In 2010s, two most remarkable regions are highlighted, i.e., north territory of Canada (e.g., Nunavut) and western states of USA. While the seasonality patterns in 1980s and 1990s are quite consistent, the changing seasonality patterns in 2000s and 2010s are truly remarkable. The shifts in seasonality patterns in the past four decades provide evidence of changing climate dynamics in North America with implications for ecological systems.
Then, we compare the seasonality patterns among different climate variables in Fig. LABEL:daymet_w12_sparsity3. The yearly seasonality patterns of precipitation demonstrate remarkable variations in the past four decades even among nearby small regions. The seasonality patterns of minimum and maximum temperature in 1980s and 1990s are quite consistent. However, the seasonality pattern of the minimum temperature in 2000s shows more seasonal regions than the maximum temperature. In 2010s, the maximum temperature of Mexico is less seasonal than the minimum temperature. the maximum temperature of western regions of USA and central regions of North America is more seasonal than the minimum temperature. In contrast, the maximum temperature of northern regions of Canada is less seasonal than the minimum temperature.
7.2.3 Sensitivity Analysis of Phase Segmentation
As mentioned above, we discover the yearly seasonality from monthly climate variable time series with each phase corresponding to each decade. In what follows, we evaluate the model with a 5-year phase on the minimum temperature data from 2000 to 2019. Fig. LABEL:fast_mip_daymet_prcp_5_year_phase_0 shows the seasonality pattern that is more consistent with Fig. LABEL:fast_mip_daymet_tmin for 1990s. Fig. LABEL:fast_mip_daymet_prcp_5_year_phase_1 depicts the seasonality pattern from 2005 to 2009 that is dominant in 2000s, see Fig. LABEL:fast_mip_daymet_tmin. Observing the seasonality pattern in Fig. LABEL:fast_mip_daymet_prcp_5_year_phase_2, more regions around the central part of North America show less seasonal minimum temperature from 2010 to 2014, comparing with the seasonality pattern in the next 5 years as shown in Fig. LABEL:fast_mip_daymet_prcp_5_year_phase_3. Notably, as the minimum temperature of Mexico from 2015 to 2019 is more seasonal than the periods from 2000 to 2014, it is difficult to see such an evidence from Fig. LABEL:daymet_w12_sparsity3.
@@ 7.3 Seasonality Patterns of Sea Surface Temperature
In this work, we consider the sea surface temperature dataset that covers monthly averages of temperatures on the spatial resolution of a (0.25 degree latitude, 0.25 degree longitude)-grid, and there are $720\times 1440$ global grids (i.e., 1,036,800 cells) in total. [^2] https://www.ncei.noaa.gov/data/sea-surface-temperature-optimum-interpolation/v2.1/access/avhrr/ The dataset spans a past four-decade period from January 1982 to December 2019. Fig. 16 shows both monthly and yearly temperature trends of the global sea surface temperature, demonstrating a gradual warming trend as evident from the increase in yearly average temperatures. These results highlight the increasing influence of climate change on ocean temperatures. The higher fluctuations in the monthly time series reveal seasonal variations. Thus, quantifying yearly seasonality is important for measuring the climate change and understanding the dynamical system because stronger seasonality of temperatures implies less variations in the time series.
@@
In the experiment, we first set the order and the sparse level as $d=12$ and $\tau=3$, respectively. We examine the STV-SAR model on the sea surface temperature data of different decades independently and compare the coefficients of yearly seasonality of temperatures. Fig. LABEL:w12_sparsity3 explores the yearly periodicity of sea surface temperature through a spatial representation of the coefficients associated with yearly seasonality. These results show the intensity of yearly sea surface temperature variations, with higher coefficients indicating stronger periodicity. High values, marked in red, are concentrated in regions close to Eastern/Middle Asia, North America, Europe, and North Africa, where seasonal temperature changes are more pronounced. In contrast, the equatorial regions such as the area of El Niño exhibit lower coefficients, demonstrating less seasonal and predictable sea surface temperature. Although there is a consistent pattern of strong periodicity across the decades, the minor shifts indicating changes in oceanic dynamics can also be identified. For instance, as shown in Fig. LABEL:w12_sparsity3_canada, these regions close to Canada have temperature variations and seasonality shift in the past three decades. One can see from the coefficients that the temperature of Arctic Ocean became less seasonal in the past two decades, potentially influenced by factors such as oceanic circulation changes, global warming, or variability in atmospheric patterns.
@@ 8 Conclusion
In this work, we provide a unified SAR framework for quantifying time series periodicity in real-world spatiotemporal systems. We advance the development of TV-SAR and STV-SAR models from an interpretable machine learning perspective by incorporating spatially- and time-varying sparsity constraints. To address the MIO problems, we present efficient and scalable algorithmic implementation by designing a DVP strategy and a two-stage optimization scheme for TV-SAR and STV-SAR, respectively. Our findings discovered by the proposed models from ridesharing trip data and climate variable data empirically demonstrate the interpretable components related to periodicity and seasonality of time series. The spatial and temporal dynamics and patterns revealed by time series periodicity also justify remarkable real-world implications of periodicity quantification to spatiotemporal systems.
Appendix A Proof of Theorem 1
Proof.
Let $\mathcal{F}_{\tau}=\{\boldsymbol{w}\in\mathbb{R}^{d}:\|\boldsymbol{w}\|_{0}\leq\tau\}$ be the class of $\tau$-sparse vectors. Since $\hat{\boldsymbol{w}}$ minimizes the residual error over $\mathcal{F}_{\tau}$, we have:
$\|\tilde{\boldsymbol{x}}-\boldsymbol{A}\hat{\boldsymbol{w}}\|_{2}^{2}\leq\|\tilde{\boldsymbol{x}}-\boldsymbol{A}\boldsymbol{w}^{\star}\|_{2}^{2}.$
Substituting $\tilde{\boldsymbol{x}}=\boldsymbol{A}\boldsymbol{w}^{\star}+\boldsymbol{\epsilon}$, this becomes:
$\|\boldsymbol{A}(\hat{\boldsymbol{w}}-\boldsymbol{w}^{\star})-\boldsymbol{\epsilon}\|_{2}^{2}\leq\|\boldsymbol{\epsilon}\|_{2}^{2}.$
Expanding the left-hand side:
$\|\boldsymbol{A}(\hat{\boldsymbol{w}}-\boldsymbol{w}^{\star})\|_{2}^{2}-2\langle\boldsymbol{\epsilon},\boldsymbol{A}(\hat{\boldsymbol{w}}-\boldsymbol{w}^{\star})\rangle+\|\boldsymbol{\epsilon}\|_{2}^{2}\leq\|\boldsymbol{\epsilon}\|_{2}^{2}.$
Cancelling $\|\boldsymbol{\epsilon}\|_{2}^{2}$ from both sides:
$\|\boldsymbol{A}(\hat{\boldsymbol{w}}-\boldsymbol{w}^{\star})\|_{2}^{2}\leq 2\langle\boldsymbol{\epsilon},\boldsymbol{A}(\hat{\boldsymbol{w}}-\boldsymbol{w}^{\star})\rangle.$
Now take expectations on both sides:
$\mathbb{E}\left[\|\boldsymbol{A}(\hat{\boldsymbol{w}}-\boldsymbol{w}^{\star})\|_{2}^{2}\right]\leq 2\cdot\mathbb{E}\left[\sup_{\boldsymbol{w}\in\mathcal{F}_{\tau}}\langle\boldsymbol{\epsilon},\boldsymbol{A}(\boldsymbol{w}-\boldsymbol{w}^{\star})\rangle\right].$
Let $\mathcal{S}_{\tau}=\{\boldsymbol{A}(\boldsymbol{w}-\boldsymbol{w}^{\star}):\boldsymbol{w}\in\mathcal{F}_{\tau}\}$. We now apply a Gaussian complexity bound:
$\mathbb{E}\left[\sup_{\boldsymbol{v}\in\mathcal{S}_{\tau}}\langle\boldsymbol{\epsilon},\boldsymbol{v}\rangle\right]\leq\sigma\cdot\sqrt{2\log\left(|\mathcal{F}_{\tau}|\right)}\cdot\sup_{\boldsymbol{v}\in\mathcal{S}_{\tau}}\|\boldsymbol{v}\|_{2}.$
The number of $\tau$-sparse supports is bounded by:
$|\mathcal{F}_{\tau}|\leq\binom{d}{\tau}\leq\left(\frac{ed}{\tau}\right)^{\tau}.$
Taking logarithms:
$\log|\mathcal{F}_{\tau}|\leq\tau\log\left(\frac{ed}{\tau}\right)\leq\tau\left(\log\left(\frac{d}{\tau}\right)+1\right).$
Putting everything together:
$\mathbb{E}\left[\|\boldsymbol{A}(\hat{\boldsymbol{w}}-\boldsymbol{w}^{\star})\|_{2}^{2}\right]\leq C\cdot\sigma^{2}\cdot\tau\left(\log\left(\frac{d}{\tau}\right)+1\right),$
for some universal constant $C$. Finally, divide both sides by $T-d$ to normalize by the number of samples:
$\mathbb{E}\left[\frac{1}{T-d}\|\boldsymbol{A}(\hat{\boldsymbol{w}}-\boldsymbol{w}^{\star})\|_{2}^{2}\right]\leq C\cdot\sigma^{2}\cdot\frac{\tau}{T-d}\left(\log\left(\frac{d}{\tau}\right)+1\right).$

### v2

@@ 7.1.2 Model Training: Comparison of Estimation Algorithms
This highlights an important tradeoff: although MIO guarantees the best solution, MIO-DVP offers a scalable alternative with much faster runtimes due to its reduced search space. The pruning parameter $\tau_{0}$ plays a critical role—aggressive pruning can degrade performance, whereas moderate values (like $\tau_{0}=10$) enable a favorable balance between efficiency and accuracy.
7.1.3 Results: Periodicity of Ridesharing Trips
To capture the temporally-varying dynamics present in the dataset, we apply the proposed TV-SAR model, treating each month as an independent time segment. The autoregressive order is set to $d=168$, corresponding to one week of hourly lags. We use a sparsity level of $\tau=4$ to allow the model to select dominant lags, particularly targeting daily ($k=24$) and weekly ($k=168$) periodicities.
Figure LABEL:NYC_JFK_w_tau_3 shows that the model consistently recovers the support set $\operatorname{supp}(\boldsymbol{w}_{\gamma})={1,24,167,168}$ for all $\gamma\in[2,60]$, reflecting the presence of strong local, daily, and weekly auto-correlations. Figure LABEL:NYC_JFK_pickup_w_tau_3 visualizes how the periodic structure of pickup trips evolves over time. In 2020, the strength of the weekly component decreases remarkably, while the daily component grows stronger. This pattern is similarly observed in the dropoff coefficients (Figure LABEL:NYC_JFK_dropoff_w_tau_3), where we also note a sharp increase in the local coefficient at $k=1$ during 2020—suggesting increased short-term variability in dropoff behavior.
Looking across all years, the weekly periodicity at $k=168$ is generally more pronounced in dropoff trips than in pickup trips. This is consistent with the patterns seen in Figure LABEL:NYC_JFK_airport_heatmap, where dropoff activity follows more regular flight schedules, while pickups are affected by greater uncertainty (e.g., baggage claim, rideshare wait times). Although periodicity in 2020 differs substantially from the adjacent years, the autoregressive patterns in 2019, 2021, 2022, and 2023 show no significant differences. This indicates that by 2021–2023, human mobility behavior had largely returned to its pre-pandemic structure.
7.2 Seasonality Patterns of North America Climate Variables
This case study investigates large-scale seasonal patterns in climate variables across North America using long-term temperature and precipitation data (1980–2019). These data exhibit strong yearly seasonality influenced by geography and climate change, and are structured both spatially and temporally. Our goal is to quantify the strength and variability of these seasonal patterns over space and time. We use the STV-SAR model to uncover the geographical organization of seasonality, track its evolution over decades, and assess the robustness of the learned patterns across spatial resolutions and segmentation schemes.
7.2.1 Dataset and Motivation for STV-SAR
We study the Daymet dataset, which provides monthly climate variables—including minimum/maximum temperature and precipitation—at a high spatial resolution of 1km×$\times$1km across North America [34]. [^1] https://daac.ornl.gov/DAYMET Our analysis focuses on data from 1980 to 2019, resulting in long, spatially-distributed multivariate time series. Figure 10 illustrates the minimum temperature data for January over the past decade. These panels reveal strong seasonal structure and regional climate variability, suggesting the need for models that can capture recurring periodic patterns as well as their changes across space and time.
@@ 7.2.1 Dataset and Motivation for STV-SAR
Given the spatial scale and long-term coverage of this dataset, we aim to (i) quantify yearly seasonality and its regional variation, and (ii) examine how such patterns evolve across decades and spatial resolutions. These goals motivate the use of our proposed STV-SAR model, which is well-suited for discovering global periodic structures in large-scale, multidimensional time series.
7.2.2 Global Support Verification
To enable scalable model training on high-resolution spatiotemporal data, we aggregate the original 1 km×$\times$1 km dataset to a coarser resolution of 10 km×$\times$10 km. Even at this level, the STV-SAR model of Eq. (9) must be fit across over 800,000 time series. To manage this computational burden, we employ the two-stage optimization scheme outlined in Sections 6.2 and 6.3. As shown in Table III, our scheme trains the full model across all time series in under 12 seconds—demonstrating its practical scalability for large-scale climate datasets.
We apply STV-SAR with a prescribed sparsity level $\tau=3$, aiming to capture dominant temporal dependencies that characterize yearly seasonality. For example, using the minimum temperature dataset, we construct monthly time series at each spatial cell over decadal segments (from 1980s to 2010s). Solving the global model yields a common support set $\Omega=\{1,11,12\}$, which includes recent memory ($k=1$), near-annual lags ($k=11,12$), and aligns with expected seasonal dynamics.
To assess the robustness of this global support set, we also fit independent SAR models to 204,153 time series at 10 km×$\times$10 km resolution using sparsity levels $\tau\in\{3,4,5\}$. As shown in Figure LABEL:daymet_supp_set, even without enforcing any global constraint, the most frequently selected lags coincide with the global support set $\Omega$. For $\tau=3$, the dominant indices are $\{1,11,12\}$; for higher sparsity levels, the support expands to include nearby lags (e.g., $\{2,10\}$), but $\{1,11,12\}$ remain consistently prominent. This confirms that the global support set not only enhances computational efficiency but also meaningfully captures shared temporal structures across space.
7.2.3 Results: Seasonality Patterns across Four Decades
Understanding how seasonality patterns evolve over time allows us to quantify the dynamics of large-scale climate systems. Figure LABEL:daymet_w12_sparsity3 visualizes the yearly seasonality of monthly climate time series across North America from 1980 to 2019, using a spatial resolution of 10 km×$\times$10 km. Higher coefficients correspond to regions with stronger yearly periodicity.
We begin by analyzing the seasonality patterns of minimum and maximum temperature. Across the four decades, areas with stronger yearly seasonality are consistently concentrated in high-latitude regions, such as Canada and Alaska. In contrast, southern regions like Mexico exhibit weaker and more spatially variable seasonality. A notable expansion of highly seasonal regions occurs in the 2000s, suggesting an intensification of temperature-driven seasonal cycles. In the 2010s, two regions stand out: northern Canada (e.g., Nunavut) and the western United States. While the patterns in the 1980s and 1990s are relatively stable, the 2000s and 2010s show substantial shifts—highlighting evolving climate dynamics with potential ecological implications.
Next, we compare seasonal structures across different climate variables. The seasonality of precipitation exhibits significant spatial heterogeneity throughout the decades, even across small neighboring regions. Meanwhile, the seasonality of minimum and maximum temperatures remains relatively stable in the 1980s and 1990s. However, in the 2000s, minimum temperature seasonality intensifies and expands more broadly than that of maximum temperature. In the 2010s, Mexico shows higher seasonality in minimum temperature compared to maximum temperature, whereas the reverse holds in the western U.S. and central North America. Notably, in northern Canada, maximum temperature is less seasonal than minimum temperature. These contrasting patterns suggest differing drivers and sensitivities of seasonal cycles across variables and geographies.
7.2.4 Sensitivity Analysis: Spatial Resolution
We now investigate the effect of spatial aggregation by comparing results at three different resolutions: 5 km×$\times$5 km, 10 km×$\times$10 km, and 20 km×$\times$20 km. Figure LABEL:fast_mip_daymet_20km_tmin_w12_sparsity3_10s illustrates that the spatial distribution of yearly seasonality remains consistent across these resolutions. In particular, the northern regions of Canada, the western United States, and the southeastern United States exhibit strong seasonal signals, while central North America displays substantially weaker seasonality.
Temperature seasonality in Mexico shows high spatial variability and inconsistency among nearby areas, as highlighted in Figure LABEL:fast_mip_daymet_mexico_10s. Fine-grained resolution (e.g., 5 km) reveals localized structure that may be smoothed out at coarser scales. This demonstrates that our STV-SAR model captures robust patterns across resolutions while still preserving high-resolution detail where available.
To address the computational challenges of large-scale training, we employ a two-stage optimization scheme that efficiently solves the STV-SAR problem across millions of time series. Table III reports runtime statistics at each resolution. Remarkably, the model computes over 3 million time series at 5 km×$\times$5 km resolution within one minute, underscoring the scalability of the proposed method for real-world geospatial applications.
@@ 7.2.4 Sensitivity Analysis: Spatial Resolution
TABLE III: Runtime of STV-SAR on the North America (minimum) temperature dataset across four decades with different spatial resolutions. Note that $\nu$ denotes the number of time series.
5 km×$\times$5 km 10 km×$\times$10 km 20 km×$\times$20 km
@@
Runtime (s) 48.55 11.77 2.88
7.2.5 Sensitivity Analysis: Phase Segmentation
In our main analysis, we segmented the climate time series into decades to capture long-term trends. We now evaluate the sensitivity of our findings to the choice of segmentation granularity by re-estimating seasonality patterns using shorter, 5-year phases on the minimum temperature data from 2000 to 2019. Figure LABEL:fast_mip_daymet_prcp_5_year_phase_0 illustrates the spatial distribution of yearly seasonality from 2000 to 2004, which closely resembles the 1990s pattern shown in Figure LABEL:fast_mip_daymet_tmin. In the 2005–2009 window (Figure LABEL:fast_mip_daymet_prcp_5_year_phase_1), seasonality intensifies in several regions, aligning with the dominant pattern of the 2000s. From 2010 to 2014 (Figure LABEL:fast_mip_daymet_prcp_5_year_phase_2), central North America shows noticeably weaker seasonal signals, whereas the period from 2015 to 2019 (Figure LABEL:fast_mip_daymet_prcp_5_year_phase_3) marks a rebound in seasonality strength, particularly in Mexico. These subtle temporal shifts are less evident in the decade-level summary (Figure LABEL:daymet_w12_sparsity3), highlighting the value of finer segmentation for detecting short-term variations in seasonality patterns.
@@ 7.3 Seasonality Patterns of Sea Surface Temperature
This case study focuses on identifying and tracking yearly periodic patterns in global sea surface temperatures from 1982 to 2019. Sea surface temperature (SST) is a key indicator of climate variability, with dynamics influenced by geography, ocean circulation, and anthropogenic warming. The data exhibit both strong seasonality and long-term trends, making them ideal for evaluating the ability of STV-SAR to capture structured temporal variation at scale. We use the model to uncover geographic patterns of seasonal strength and examine how these patterns evolve across decades and spatial regions.
7.3.1 Dataset and Motivation for STV-SAR
We study the NOAA Optimum Interpolation Sea Surface Temperature (OISST) dataset, which provides monthly average temperatures over a $0.25^{\circ}\times 0.25^{\circ}$ global grid, comprising $720\times 1440$ cells (i.e., 1,036,800 spatial locations). [^2] https://www.ncei.noaa.gov/data/sea-surface-temperature-optimum-interpolation/v2.1/access/avhrr/ The dataset spans four decades, from January 1982 to December 2019.
Figure 14 presents both monthly and yearly average sea surface temperatures across this period. The monthly time series reveals strong seasonal fluctuations, while the yearly average trend shows a gradual warming signal consistent with global climate change. These dynamics highlight the importance of quantifying yearly seasonality—not only to understand recurring oceanic patterns but also to characterize the variability and potential instability in the climate system. In particular, stronger seasonality implies more predictable, cyclic dynamics, whereas weaker seasonality can indicate structural shifts or increased variability.
@@ 7.3.1 Dataset and Motivation for STV-SAR
7.3.2 Results: Spatial and Temporal Structure of Sea Surface Seasonality
To analyze the temporal structure of sea surface temperatures, we fit the STV-SAR model separately on each decade of data. We fix the model order at $d=12$ and sparsity level $\tau=3$, which allows us to isolate and interpret coefficients corresponding to yearly periodicity.
Figure LABEL:w12_sparsity3 shows the spatial distribution of the coefficient corresponding to lag 12, which serves as a proxy for yearly seasonality strength. Higher values (in red) indicate regions with stronger temperature seasonality—most notably near Eastern and Central Asia, North America, Europe, and North Africa. In contrast, equatorial regions such as the Pacific Ocean around the El Niño zone exhibit lower coefficients, reflecting weaker seasonality and more irregular variation.
Despite the broad stability of global seasonal patterns across decades, we observe subtle temporal changes. For example, as shown in Figure LABEL:w12_sparsity3_canada, the regions near Canada and the Arctic Ocean exhibit decreasing seasonality over time. This decline may signal evolving oceanic dynamics due to changes in ocean circulation, warming trends, or atmospheric variability. The Arctic Ocean, in particular, appears to have become notably less seasonal over the past two decades, potentially reflecting broader climate-induced disruptions in polar ocean systems.
@@ 8 Conclusion
This work introduces a unified SAR framework for quantifying periodicity in real-world spatiotemporal time series. We develop interpretable extensions—TV-SAR and STV-SAR—that incorporate temporally and spatially-varying sparsity constraints to reveal structured periodic components. To address the computational challenges posed by the underlying MIO problems, we propose scalable algorithmic solutions: a DVP-based method for TV-SAR and a two-stage optimization scheme for STV-SAR. Applied to both ridesharing mobility data and climate time series, our models successfully uncover meaningful, interpretable seasonal patterns and periodic structures. The learned spatial and temporal dynamics illustrate the practical relevance of periodicity quantification for analyzing complex systems.
We point out the following directions for future improvement. The global support estimation strategy used in STV-SAR is particularly effective when temporal structures are shared across time series, as in climate datasets (e.g., Figure LABEL:daymet_supp_set). However, for heterogeneous datasets with diverse or complex dynamics, the approximation introduced by the two-stage optimization may lead to estimation bias. Addressing this gap—potentially by integrating adaptive or hierarchical modeling techniques—could further enhance the robustness and generality of the framework.